OpenStudy (anonymous):

The area of a triangle is 63 square inches. The base is 18 inches long. What is the height of the triangle? A. 3.5 in. B. 7 in. C. 14 in. D. 31.5 in.

OpenStudy (anonymous):

Area of a triangle = 1/2 * base *height 63 =1/2 * 18*height height = 63 /9 height =7 in